# Towards an Explainable Comparison and Alignment of Feature Embeddings

### v2

@@ A.3 Proof of Theorem 2
The last line in the above inequalities follow from Young’s inequality showing that $\|A+B\|_{F}^{2}\leq 2\|A\|_{F}^{2}+2\|B\|_{F}^{2}$. Then, setting $\delta=2\exp\bigl{(}\frac{8-m\epsilon^{2}}{32}\bigr{)}$ implying $\epsilon=\sqrt{\frac{32\log(2e^{-1/4}/\delta)}{m}}$, we will have
[…]
@@
$\mathbb{P}\Bigl{(}\Bigl{\|}\frac{1}{n}\widehat{K}_{\psi_{1}}-\frac{1}{n}K_{\psi_{1}}\Bigr{\|}_{F}\leq\sqrt{\frac{32\log(2e^{-1/4}/\delta)}{m}}\;\;\text{and}\;\;\Bigl{\|}\frac{1}{n}\widehat{K}_{\psi_{2}}-\frac{1}{n}K_{\psi_{2}}\Bigr{\|}_{F}\leq\sqrt{\frac{32\log(2e^{-1/4}/\delta)}{m}}\Bigr{)}\,\geq\,1-\delta$
@@
[…]
The above completes the theorem’s proof.
A.4 Proof of Proposition 2
@@ Appendix B Additional Numerical Results
B.1 Ablation Study on the Comparison of Visualization Algorithms
@@
B.2 Comparison of text embeddings
Synthetic text dataset. To evaluate the performance of SPEC on text embeddings, we generated a dataset of 10K text samples using GPT-4o, covering diverse categories such as profession, objects, gender, and emotions. We then applied SPEC to compare CLIP, RoBERTa, and E5 text embeddings. As shown in Figures 8 and 9, CLIP primarily clusters sentences based on objects mentioned in the text, whereas RoBERTa organizes them according to profession and gender. For instance, CLIP’s first cluster consists of sentences related to cameras and photography, while RoBERTa’s first cluster groups sentences about female firefighters and female carpenters. This suggests that CLIP embeddings do not cluster professions based on gender but excel at grouping objects, which aligns with its training focus. Additionally, the t-SNE visualization reveals that the principal clusters identified in one embedding are not well-separated in the other, indicating that each model captures different semantic aspects of the text. In addition to previous experiments, we used MS-COCO 2017 train set captions ( 120K samples) to compare RoBERTa and E5-L-V2 embeddings. As shown in Figure 10, we can observe that E5 managed to cluster captions where two or more animals are interacting.
@@ B.2 Comparison of text embeddings
B.3 Comparison of image embeddings
@@
B.4 Comparing similarity ranking for SPEC clusters
In Figure  12, the leftmost images show the top 4 samples of SPEC-identified clusters on AFHQ. The first and second clusters correspond to black cats and black & white cats, respectively. For each cluster, cosine similarity is computed between the mean of these samples and both 4 additional cluster members (green-bordered) and 4 random images (red-bordered). Images are sorted left to right by similarity scores. The first row represents DINOv2, the second CLIP. Unlike DINOv2, CLIP ranks some random cats more similar to the cluster mean than the actual cluster members.
For the FFHQ dataset, as shown in Figure  13, the first cluster corresponds to people wearing graduation caps, and the second to people wearing sunglasses. While DINOv2 maintains a significant similarity gap, clearly distinguishing the clusters, CLIP assigns higher similarity to some samples that lack these defining features.
@@ B.4 Comparing similarity ranking for SPEC clusters
B.5 SPEC-align Experiments
SPEC-align Finetuning Experiments Parameters. The following parameters were used in our experiment. We used the OpenCLIP GitHub repository (link) and used the MS-COCO 2017 training set which consists of  120K pairs of texts and images. We use SPEC-align with the following parameters and chose DINOv2-Vit-B/14 as our reference model.

### v3

@@ A.4 Proof of Theorem 2
The last line in the above inequalities follow from Young’s inequality showing that $\|A+B\|_{F}^{2}\leq 2\|A\|_{F}^{2}+2\|B\|_{F}^{2}$. Then, setting $\delta=2\exp\bigl{(}\frac{8-m\epsilon^{2}}{32}\bigr{)}$ implying $\epsilon=\sqrt{\frac{32\log(2e^{1/4}/\delta)}{m}}$, we will have
$\mathbb{P}\Bigl{(}\Bigl{\|}\frac{1}{n}\widehat{K}_{\psi_{1}}-\frac{1}{n}K_{\psi_{1}}\Bigr{\|}_{F}\leq\sqrt{\frac{32\log(2e^{1/4}/\delta)}{m}}\Bigr{)}\,\geq\,1-\frac{\delta}{2},\quad\mathbb{P}\Bigl{(}\Bigl{\|}\frac{1}{n}\widehat{K}_{\psi_{2}}-\frac{1}{n}K_{\psi_{2}}\Bigr{\|}_{F}\leq\sqrt{\frac{32\log(2e^{1/4}/\delta)}{m}}\Bigr{)}\,\geq\,1-\frac{\delta}{2}$
@@
$\mathbb{P}\Bigl{(}\Bigl{\|}\frac{1}{n}\widehat{K}_{\psi_{1}}-\frac{1}{n}K_{\psi_{1}}\Bigr{\|}_{F}\leq\sqrt{\frac{32\log(2e^{1/4}/\delta)}{m}}\;\;\text{and}\;\;\Bigl{\|}\frac{1}{n}\widehat{K}_{\psi_{2}}-\frac{1}{n}K_{\psi_{2}}\Bigr{\|}_{F}\leq\sqrt{\frac{32\log(2e^{1/4}/\delta)}{m}}\Bigr{)}\,\geq\,1-\delta$
@@
$\mathbb{P}\Bigl{(}\sum_{i=1}^{n}\Bigl{\|}\Lambda_{\psi_{1},\psi_{2}}\widehat{\mathbf{v}}_{i}-{\lambda}_{i}\widehat{\mathbf{v}}_{i}\Bigr{\|}^{2}_{2}\leq{\frac{128\log(2e^{1/4}/\delta)}{m}}\Bigr{)}\,\geq\,1-\delta$
The above completes the theorem’s proof given that $2e^{1/4}\approx 2.57<3$.
A.5 Proof of Proposition 2
@@ Appendix B Additional Numerical Results
B.1 Cholesky decomposition-based method for eigenvectors and eigenvalues computation
As explained in Remark 1, the computation of the eigenvalues and eigenvectors of the differential covariance matrix $\Gamma_{\psi_{1},\psi_{2}}$ can be efficiently reduced to the eigendecomposition of a symmetric matrix through the use of Cholesky decomposition. We followed the explanation in the Appendix A.3 and compared the runtime between Algorithm 1 and the Cholesky decomposition-based method on the Gaussian Kernel with different RFF. As shown in Table 1, the Cholesky method is more scalable and was able to handle $d=10K$, whereas the vanilla implementation of Algorithm 1 using Pytorch’s eig command (for general square matrices) could not efficiently scale beyond dimension $d=5000$.
Furthermore, in Figures 6, 7, we present the top four SPEC-identified clusters comparing CLIP with DINOv2 using different values of Random Fourier Features (RFF) $r$, ranging from $1000$ to $6000$, with Cholesky decomposition applied to the MS-COCO dataset. We observe that for $r\leq 3000$, the clusters are not well-separated and show some inconsistencies. As $r$ increases, these inconsistencies diminish, and by $r=6000$, the clusters are clean and well-separated. We also provide the t-SNE and UMAP representations of the top 10 SPEC-identified clusters for CLIP and DINOv2 for r=2000, 6000r=2000,\,6000. in Figures 8, 9.
Time (sec)
Computation Method d=d=1000 d=d=2000 d=d=4000 d=d=6000 d=d=8000 d=d=10000
Direct (Algorithm 1) 13 35 72 - - -
Cholesky Decomposition-based (Appendix A.3) 11 22 33 47 64 85
Table 1: Time taken for Direct method and Cholesky Decomposition method at different values of $d$.
Figure 6: Top 4 SPEC-identified clusters comparing CLIP - DINOv2 with different Random Fourier features (RFF) r ranging from 1000 to 6000 using Cholesky decomposition on MS-COCO dataset.
Figure 7: Top 4 SPEC-identified clusters comparing DINOv2 - CLIP with different Random Fourier features (RFF) r ranging from 1000 to 6000 using Cholesky decomposition on MS-COCO dataset.
Figure 8: Top 4 SPEC-identified clusters comparing CLIP - DINOv2 embeddings on MS-COCO dataset using Cholesky decomposition with different Random Fourier Features (2000 and 6000). The second row shows the t-SNE and UMAP representation of the top 10 SPEC-identified clusters for each embedding.
Figure 9: Top 4 SPEC-identified clusters comparing DINOv2 - CLIP embeddings on MS-COCO dataset using Cholesky decomposition with different Random Fourier Features (2000 and 6000). The second row shows the t-SNE and UMAP representation of the top 10 SPEC-identified clusters for each embedding.
B.2 Ablation Study on the Comparison of Visualization Algorithms
@@
B.3 Comparison of text embeddings
Synthetic text dataset. To evaluate the performance of SPEC on text embeddings, we generated a dataset of 10K text samples using GPT-4o, covering diverse categories such as profession, objects, gender, and emotions. We then applied SPEC to compare CLIP, RoBERTa, and E5 text embeddings. As shown in Figures 12 and 13, CLIP primarily clusters sentences based on objects mentioned in the text, whereas RoBERTa organizes them according to profession and gender. For instance, CLIP’s first cluster consists of sentences related to cameras and photography, while RoBERTa’s first cluster groups sentences about female firefighters and female carpenters. This suggests that CLIP embeddings do not cluster professions based on gender but excel at grouping objects, which aligns with its training focus. Additionally, the t-SNE visualization reveals that the principal clusters identified in one embedding are not well-separated in the other, indicating that each model captures different semantic aspects of the text. In addition to previous experiments, we used MS-COCO 2017 train set captions ( 120K samples) to compare RoBERTa and E5-L-V2 embeddings. As shown in Figure 14, we can observe that E5 managed to cluster captions where two or more animals are interacting.
@@ B.3 Comparison of text embeddings
B.4 Comparison of image embeddings
@@
B.5 Comparing similarity ranking for SPEC clusters
In Figure  16, the leftmost images show the top 4 samples of SPEC-identified clusters on AFHQ. The first and second clusters correspond to black cats and black & white cats, respectively. For each cluster, cosine similarity is computed between the mean of these samples and both 4 additional cluster members (green-bordered) and 4 random images (red-bordered). Images are sorted left to right by similarity scores. The first row represents DINOv2, the second CLIP. Unlike DINOv2, CLIP ranks some random cats more similar to the cluster mean than the actual cluster members.
For the FFHQ dataset, as shown in Figure  17, the first cluster corresponds to people wearing graduation caps, and the second to people wearing sunglasses. While DINOv2 maintains a significant similarity gap, clearly distinguishing the clusters, CLIP assigns higher similarity to some samples that lack these defining features.
@@ B.5 Comparing similarity ranking for SPEC clusters
B.6 SPEC-align Experiments
SPEC-align using $\ell_{2}$-Norm of Kernel Difference Matrix. As mentioned in Remark 2, another choice of SPEC-diff uses the $\ell_{2}$-norm of the eigenvalues of the kernel difference matrix $\Lambda_{\psi_{1},\psi_{2}}$. Similar to the experiment in Figure 5, we aligned CLIP to DINOv2 using this $\ell_{2}$-norm-based objective. Figure 22 shows the CLIP kernel before and after alignment, alongside the target DINOv2 kernel. We also plot the SPEC-diff score over iterations, demonstrating successful convergence of the alignment process.
Figure 22: SPEC-align using the $\ell_{2}$-norm of the eigenvalues of the kernel difference matrix. Left: original CLIP kernel. Center: aligned CLIP kernel using $\ell_{2}$-norm-based SPEC-align. Right: target DINOv2 kernel. Bottom: SPEC-diff score ($\ell_{2}$-Norm) over alignment iterations.
SPEC-align Finetuning Experiments Parameters. The following parameters were used in our experiment. We used the OpenCLIP GitHub repository (link) and used the MS-COCO 2017 training set, which consists of  120K pairs of texts and images. We use SPEC-align with the following parameters and chose DINOv2-Vit-B/14 as our reference model.